OpenStudy (whalenjacob):

Consider the following set of equations: Equation R: −3y = −3x – 9 Equation S: y = x + 3 Which of the following best describes the solution to the given set of equations? No solution One solution Infinite solutions Two solutions

OpenStudy (whalenjacob):

@DullJackel09

OpenStudy (whalenjacob):

@andretorres

OpenStudy (whalenjacob):

i will medal and fan sombody help

jhonyy9 (jhonyy9):

-3y = -3x -9 y=x+3 in case of the first one equation do you can factorizing out on the both sides the -3 ? so what will get like result ?

jhonyy9 (jhonyy9):

wolf do you agree this like a first step ?

jhonyy9 (jhonyy9):

jacob please collaborate

OpenStudy (whalenjacob):

sorry

jhonyy9 (jhonyy9):

do you can factorizing out the -3 ?

OpenStudy (whalenjacob):

i dont know i, so cumfused with this

OpenStudy (wolf1728):

Multiply equation S by -3

OpenStudy (wolf1728):

That will give you the answer

OpenStudy (whalenjacob):

thank you so much

jhonyy9 (jhonyy9):

yes this is indifferent - multiplie on the both sides by -3

OpenStudy (whalenjacob):

i have one more

OpenStudy (whalenjacob):

Which description best describes the solution to the following system of equations? y = −x + 4 y = 3x + 3 Line y = −x + 4 intersects the line y = 3x + 3. Lines y = −x + 4 and y = 3x + 3 intersect the x-axis. Lines y = −x + 4 and y = 3x + 3 intersect the y-axis. Line y = −x + 4 intersects the origin.

OpenStudy (wolf1728):

If the equations do have a solution, then one line will intersect the other line.

OpenStudy (wolf1728):

So I'd say Line y = −x + 4 intersects the line y = 3x + 3 would best describe the solution

OpenStudy (whalenjacob):

thx

OpenStudy (wolf1728):

u r welcome by the way the 2 equations do have a solution.